### v1

# Studying the Effects of Collaboration in Interactive Theme Discovery Systems

Abstract
NLP-assisted solutions have gained considerable traction to support qualitative data analysis. However, there does not exist a unified evaluation framework that can account for the many different settings in which qualitative researchers may employ them. In this paper, we take a first step in this direction by proposing an evaluation framework to study the way in which different tools may result in different outcomes depending on the collaboration strategy employed. Specifically, we study the impact of synchronous vs. asynchronous collaboration using two different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
11footnotetext: Equal Contribution
1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes untenable as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al. (2019).
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to automatically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al. (2018); Fang et al. (2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al. (2023), or providing live recommendations to users as they manually code the data Dai et al. (2023); Gao et al. (2024). To ensure the agency of the researchers, all of these systems seek to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
In most cases, these tools are evaluated to highlight the strengths and weaknesses of the proposed system. Examples include evaluating topic coherence with and without human input Fang et al. (2023) and contrasting machine-assisted code-book generation with a manual-only code-book generation Dai et al. (2023). In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we want to study how tools for inductive coding may result in different outcomes in different collaborative settings.
Collaborative data analysis occurs when two or more researchers pool their focus into a shared body of data to produce an agreed interpretation. Potential advantages of collaboration include producing more reliable results, a richer understanding of the data, and knowledge that transcends individual perspectives Flick (2014). While there are many ways in which collaboration can occur, in this work we look at two common but opposite qualitative research collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding can be a more flexible option for teams as they can work across different regions and time zones, whereas synchronous coding can combine diverse perspectives and streamline the annotation processes.
We test these settings by choosing two NLP tools for inductive coding, and performing coding experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the Fang et al. (2023) system, which uses a traditional human-in-the-loop topic modeling solution, and (2) the Pacheco et al. (2023) system, which models emerging themes with respect to their relationship to other high-level concepts, groups similar instances based on this, and leaves theme definitions and annotations to users, which are then propagated to the full dataset. This way, we can have a more representative understanding of how different tools can be affected in synchronous and asynchronous collaboration.
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we also perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting.
In summary, we make two main contributions:
  1. 1.
    We demonstrate how the collaboration setting in which qualitative research takes place could affect the outcome of solutions proposed in the NLP literature, by testing two NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts.
  2. 2.
    We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture.
We hope that our findings will inform the future development of language technologies that are in line with the expectations and workflow of qualitative researchers, as well offer insights into how the next generation of qualitative research tools should be evaluated so that researchers are able to choose the right tool for their coding requirements.
2 Related Work
Our work builds on previous literature on computational solutions for qualitative coding. The overarching goal of these systems is to partially automate the qualitative coding process by either learning user-defined themes interactively Pacheco et al. (2023); Gao et al. (2023) or by inducing topics or clusters in an interactive, semi-supervised manner Fang et al. (2023); Smith et al. (2018). A separate but related line of work exemplified by Gao et al. (2024) uses language models to generate label recommendations as the user performs the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to classify large portions of the dataset without requiring extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by literature from the field of Human-Computer Interaction (HCI)  Jiang et al. (2021); Feuston and Brubaker (2021); Chen et al. (2018). HCI research centers the user as the focus of study in order to identify where and how best to place NLP tools in interactive systems. We build on these findings by proposing a comprehensive method for evaluating a given system for qualitative coding in addition to conducting user interviews. To this end, we also center part of our experiment on the user to uncover differences in the interactive experience between synchronous and asynchronous environments.
A common finding in the HCI studies surveyed is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems were generally ad hoc, with experiments conducted variously in group settings  Choo et al. (2013); Hoque and Carenini (2016); Smith et al. (2018), on individual participants  Rietz et al. (2020), and through platforms such as MTurk  Zade et al. (2018). Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
3 Interactive Systems
We evaluate two interactive theme discovery systems that use different approaches to accomplish the same task. This way, we ensure that our analysis can account for variation in NLP techniques and annotation interfaces.
3.1 Fang et al. (2023)
This system is a HitL, query-driven topic model (QDTM) (Fang et al., 2021) that allow researchers to iteratively adjust clusters using a set of refinement operations. Broadly, the QDTM is a semi-supervised topic modeling algorithm that accepts input queries to extract relevant concept words. The extracted concept words are used in a two-phase framework based off the Hierarchical Dirichlet Process (HDP) (Teh et al., 2004) to perform topic modeling. In the first phase, the HDP is used to infer one "parent topic" for each concept along with irrelevant topics, where each topic is defined by a set of weighted words. In the second phase, the model expands the parent topic into a set of subtopics using the Gibbs sampling process. The words for each subtopic are added to the parent topic to produce a final set of topics, which are used to assign a distribution of topics to each document.
Given an initial topic model, users can add, remove, or swap the order of words within each topic, which alters the probability of topic assignment for documents that contain that specified word. Similarly, users can split or merge topics based on topic words; doing so does not guarantee that assigned documents will be reassigned accordingly due to the stochastic nature of the model. Users can also remove documents from a given topic, but cannot directly add documents in the same manner. In each iteration, users may freely select refinement operations for each topic before applying all selections at once. Users can undo selections prior to applying refinements and previous models are also saved if users wish to return to a previous iteration.
The overarching goal of this system is to facilitate the discovery of an optimal topic model for a given input dataset. By applying a selection of refinements, users can compare the effects of different topic word allocations on the resulting model. Returning to previous models and applying new refinements creates a branching model, which provides users with a wide space for exploration. Once satisfied with a given iteration, the user can download that model with its document distribution.
3.2 Pacheco et al. (2023)
This system uses a two-stage framework to assist researchers in characterizing large textual collections. In the first stage, the system automatically proposes an initial partition of the data, such that textual instances that are thematically similar are clustered together. Users are then provided with a set of exploratory operations that allows them to evaluate the quality of the discovered partitions, as well as to further explore and partition the space by inspecting individual examples, finding similar instances, and using open text queries. Next, users are prompted to group identified patterns into general themes, and instantiate them by giving them a name. Once a theme is created, users are provided with a set of intervention operations to explain the themes using natural language, select good example instances, write down additional examples, and input or correct supporting concepts to characterize the theme assignments. The full set of operations supported by the system are outlined in App. A.
In the second stage, the system finds a mapping between the full set of instances and the themes instantiated by the users. It takes the information contributed by the users in the form of examples and supporting concepts, and learns to map instances to themes using a neuro-symbolic procedure implemented with DRaiL Pacheco and Goldwasser (2021). The main goal of this process is to condition new theme assignments not only on the embedding distance between instances and provided examples, but also leverage the additional judgements provided by users via supporting concepts. For example, when analyzing a corpus about the Covid-19 vaccine, users could point out that 80% of the examples for theme Government Distrust have a clear anti-vaccine stance. The system then uses this information to introduce inductive bias into the mapping procedure. DRaiL uses weighted first-order logic rules to express decisions and dependencies between different decisions, for example $w:\mathtt{stance(anti\_vax)}\rightarrow\mathtt{theme(gov\_distrust)}$. For more details about this implementation, we refer the reader to the original paper.
During this process, the systems keeps instances unassigned to a theme if there is not a good enough match. Then, a consecutive partitioning step is done over all unassigned instances. Finally, users have the option to repeat this two-stage process until convergence.
4 Coding Protocol
In order to study the performance of the two systems discussed in Section 3 under different collaboration settings, we design a protocol to conduct qualitative coding in the synchronous and asynchronous qualitative coding paradigms. We use the outcome of our coding described here to gather results, and then conduct an evaluation in the subsequent sections.
4.1 Participants
We recruited a group of 15 researchers in NLP and Computational Social Science, 3 female and 12 male, within the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. All participants were either well versed in qualitative data analysis, or were explicitly trained by more senior researchers to perform the task.
4.2 Dataset
We use a common dataset to perform qualitative coding using the Pacheco et al. (2023) and Fang et al. (2023) systems. This dataset was originally released by Pacheco et al. (2022), and consists of approximately $85,000$ Tweets about the COVID-19 vaccine. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain predictions for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) Haidt and Graham (2007); Roy et al. (2021).
4.3 Qualitative Coding Protocol
In order to compare the consistency between different collaborative settings, we need at least two sets of synchronous/asynchronous annotations per annotation system. To ensure compatibility between synchronous and asynchronous outcomes, we use the same starting clusters for each system. The hyperparameters used in the Fang et al. (2023) system are the same as in their study and we did not use any input queries prior to initialization. We scheduled annotations for each setup separately, with details outlined in the sections below.
4.3.1 Synchronous Coding
Coding for the Pacheco et al. (2023) system was done over Zoom, where participants met online and collaboratively discussed and annotated the dataset. Six participants were bifurcated into two groups of three individuals, with each group communicating only with their members during the annotation process. Coders followed the protocol outlined in the original paper, performing two consecutive rounds of interaction, each round lasting a total of 1 hour.
For the Fang et al. (2023) system, we selected a different set of six annotators and divided them into two separate groups of three individuals each. Annotators met with their group and collaboratively coded the data. Prior to the annotation session, participants were briefed about the different features of the system and how to use them. The annotators went over the pre-assigned topics proposed by the system, and modified them to make them more accurate and cohesive. Both groups of annotators reached saturation after multiple passes over the data, in sessions lasting around 2.5 hours.
4.3.2 Asynchronous Coding
The asynchronous coding processed followed the same outline as the synchronous coding for each system, except that it was conducted by three different annotators, and it was done individually without any collaboration. For each system, the annotators first met in-person and started annotations separately. The purpose of the initial co-located coding session was to troubleshoot any problems and solve any doubts that the participants may have. The participants then completed the annotation completely asynchronously and remotely.
Pacheco et al. (2023) Fang et al. (2023)
Jaccard Sim. Centroid Sim. Group Avg. Sim. Jaccard Sim. Centroid Sim. Group Avg. Sim.
Sync $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$
Asyn $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$
Table 1: Avg. Consistency between Best Theme Matches across Annotator Groups. Statistically significant using a two-sample unpaired t-test with $p<0.05$. ∗∗ Near statistically significant with $p\approx 0.05$.
5 Evaluation
In this section, we use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. We first define and present several metrics for theme induction quality based on overlap, similarity/distance, association and verification that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al. (2021); Pacheco et al. (2023). These metrics are calculated using the sets of documents assigned to each theme by their respective system. While the Pacheco et al. (2023) system explicitly assigns a single theme for each document, the Fang et al. (2023) system provides a distribution of theme weights for each document. To ensure that results from both systems are comparable, we assign each document to the theme with the highest weight.
Finally, we summarize annotator experiences to provide a perspective on qualitative differences between the two settings. Combining dimensions in our evaluation methodology, we demonstrate that there are significant differences in consistency, cohesiveness, correctness, and user experiences between synchronous and asynchronous experiments.
5.1 Consistency
Coding individually always runs the risk that coders overgeneralize or miss key themes, resulting in “unconstrained or unsystematic” analysis (Cornish et al., 2014). We propose consistency as one dimension for measuring this systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts.
There are two obstacles to measuring annotation consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents and themes with different names may share documents that are semantically similar, which we address by measuring the consistency of documents between given pairs of themes. Following the definition of consistency in Section 4 of Pacheco et al. (2023), we calculate the Jaccard similarity given as:
$J(A,B)=\frac{|A\cap B|}{|A|+|B|-|A\cap B|}$ (1)
where $A$ and $B$ are sets of documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
Second, since the majority of documents are assigned a theme by the system, the same themes may contain a different assortment of documents but remain relatively semantically similar. We therefore also measure the semantic similarity between themes using their SBERT embeddings Reimers and Gurevych (2019). We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise similarity between all documents in the two themes. As in the Jaccard case, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings. In all cases, we use the cosine similarity.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. Comparing the results on the synchronous/asynchronous dimension, we find that synchronous groups produce themes with higher average consistency across the board. The centroid similarities and group average similarities in particular are statistically significant, supporting our hypothesis that different settings produce noticeably different coding results. These findings also demonstrate that the theoretical advantage for collaborative coding holds empirically true when using HitL systems as well: coders are more likely to miss themes or overgeneralize when working independently.
Comparing between consistency metrics, we find that centroid similarity is at or near statistical significance for all experiments and is much higher than the other metrics across the board. This finding is also in line with our expectations detailed above: since HitL systems automate a large portion of document assignments, small differences in user interactions can result in larger differences in document mappings. However, Jaccard similarity is still important to consider, especially for datasets comprised of documents that are very semantically similar.
5.2 Cohesiveness and Distinctiveness
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-Theme
Sync $\bf 0.51(0.08)^{*}$ $0.41(0.04)^{*}$ $0.52(0.10)$ $0.40(0.04)$
Async $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.52(0.10)$ $0.40(0.04)$
Table 2: Group Avg. Similarity within Annotator Groups (All data points). Themes are considered to be more cohesive if intra-theme similarity is high, and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
A second dimension for determining the "systematicity, clarity, and transparency" of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book (Cornish et al., 2014). We propose two metrics, cohesiveness and distinctiveness, to represent these similarities and differences respectively. Whereas consistency measures the systematicity and clarity of themes across annotators, cohesiveness and distinctiveness measure these qualities for themes by the same annotator. The distinctiveness of a theme is determined by comparing it to other themes created by the same annotator. On the other hand, the cohesiveness of a single theme is determined by its semantic homogeneity. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
To evaluate these metrics for a single annotator (or annotator group), we calculate both the intra-theme similarity and the inter-theme similarity for theme in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes by the same annotator.
A confounding factor in these measures is that both systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. In order to more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the Pacheco et al. (2023) system, this top quartile is selected using the distance from the centroid. For the Fang et al. (2023) system, the distribution of theme weights better reflect the assignments provided by the topic model and are used instead.
Table 2 shows results for the whole dataset and Table 3 show the same metrics for the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the Pacheco et al. (2023) experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings hint that themes may be more cohesive but not more distinctive in the synchronous setting.
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-theme
Sync $\bf 0.70(0.08)^{*}$ $0.52(0.07)^{*}$ $0.56(0.11)$ $0.39(0.05)$
Async $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.56(0.10)$ $0.39(0.05)$
Table 3: Group Avg. Similarity within Annotator Groups (25% most related). Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
The results in the Fang et al. (2023) experiments are nearly identical across collaboration modalities. Narrowing down to the top 25% of closest documents increases the difference between the two metrics, lining up with the intuitive notion that doing so increases both cohesiveness and distinctiveness. When comparing the two systems, we find that the themes coded using the Pacheco et al. (2023) system are significantly more cohesive when looking at the quartile subset, which matches our findings in Section 5.3.
5.3 Correctness
While the previous evaluations look at the code-book as a hole, we want to also evaluate how correct our document to theme assignments are. In order to verify this, we conducted a post-hoc analysis of how accurate theme assignments are on a held-out set of randomly sampled tweets, compared across synchronous and asynchronous settings. The data was sampled after the annotation process was over. This meant that for both systems, annotators had deemed the theme discovery and refinement to have reached saturation.
Following Section 5.2, we perform one analysis of correctness that considers the set of all tweets, as well as a second analysis considers only the top $25\%$ of tweets most closely related to each theme. For each analysis, we sampled $200$ tweets randomly and had two of the authors of the paper manually annotate whether the theme assigned to a tweet by the system makes sense or not. We did this for each system and collaboration setting (synchronous vs. asynchronous), with the exception of the Pacheco et al. (2023) synchronous setting, for which we report the results presented in the original paper. We made sure to follow the same sampling procedure so that results were comparable.
Results are presented in Tab. 4. We observe that for both systems, the accuracy yielded from synchronous annotations is higher than their asynchronous counter parts. We also see that as per our expectations, the top $25^{th}$ percentile has a higher accuracy than group average of all tweets.
However, it is interesting to observe the comparatively poor accuracy of the topic modeling system Fang et al. (2023) as compared to the theme discovery system Pacheco et al. (2023). We attribute this difference to the fact that topic modeling systems do not "learn" to classify themes, but rather learn the probability that a given word belongs to a particular topic. During evaluation, we found that the main problem arises when salient words are used in atypical contexts, or when trying to bifurcate terms that are used in opposing themes.
As an example, take the following headline regarding the COVID-19 vaccine mandate.
Indiana University students must comply with the school’s COVID-19 vaccine mandate for the fall semester after a federal judge refused on Monday to block the school’s inoculation requirement, one of the first rulings of its kind during the pandemic
The topic modeling system gave the highest wright to the topic Anger over mask mandates. But really, this is regarding vaccines. The term mandate biases the model to rank a topic about masks (the context that typically appears around the word "mandate") higher than a topic related to the COVID-19 vaccine, which this headline is about.
Our findings regarding the correctness can be summarized by acknowledging that synchronous teams of annotators achieve significantly higher assignment accuracy than asynchronous teams. We also find that topic modeling systems achieve un-competitive assignment accuracy versus theme discovery systems. Lastly, qualitative researchers may want to choose theme discovery tools if they want higher accuracy in document to theme assignments – but could opt for topic modeling systems if they want to put their focus on word usage patterns.
Pacheco et al. (2023) Fang et al. (2023)
All Top 25% All Top 25%
Sync ${}^{*}68\%$ ${}^{*}86\%$ $23\%$ $42\%$
Async $50\%$ $64\%$ $17\%$ $35\%$
Table 4: Correctness: Accuracy of Theme Assignment for Held-out Dataset. Results taken from Pacheco et al. (2023).
5.4 User Study
We carried out semi-structured interviews with all the participants in the study after they were done with their coding task. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. We summarize our main findings below.
Synchronous annotators found working in teams to ease the annotation process.
Both teams of annotators working synchronously using the Fang et al. (2023) described their perceived advantage of working in teams – broader perspectives by inviting view points from different annotators helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve contested tweets through discussions. Particularly, it was easy to "break ties" quickly through conversation when it came to mapping which words belonged to a given a topic. The annotators were also quicker to reach saturation, and they expressed that they thought this was possible because consolidation was possible on-the-fly. One of the annotators noted that a pre-requisite for successful synchronous coding stemmed from the rapport their annotation team shared, as they are all members of the same lab and were able to convey their viewpoints without hesitation.
Asynchronous annotators were more likely to notice inadequacies of the annotation tool they used.
While most of the highlights from the synchronous annotators focused on the process of annotation, the asynchronous annotators directed their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. The synchronous annotators also met with similar issues, but were quick to solve it collaboratively. Even though all annotators (in both settings and systems) started with the same initial clusters of tweets, the synchronous annotators were able to find the operation they need performed quicker through in-team collaboration.
Annotators noted the lack of control and agency that the topic modeling system had over the theme discovery system.
Annotators expressed loss of control and agency when using the topic modeling system Fang et al. (2023) over the theme discovery system Pacheco et al. (2023). One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups".
Annotators felt that starting topics often negatively impacted their annotations.
Since topic models start by a sampled cluster of data, both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al. (2023) model groups together despite the stark difference in message between the two.". Since the topics contain a large number of documents, it is only feasible for annotators to include / exclude a very small subset as being part of the topic. Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of clustering already done - as opposed to the Pacheco et al. (2023) system where annotations had to be done from scratch, posing to be a small hurdle in the beginning.
6 Conclusion and Future Work
In this paper, we examined NLP-assisted qualitative research tools in different collaboration modalities. We conducted annotations on a vast dataset using two popular interactive coding tools, Pacheco et al. (2023) and Fang et al. (2023). We then designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for theme discovery systems like Pacheco et al. (2023), where users have a wider bandwidth of interaction that can benefit from group consensus. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and highly accurate document-level annotations.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy in supporting qualitative coding. We believe that the evaluation framework that we proposed can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.
7 Limitations
The study presented in this paper has three main limitations.
(1) While we selected two distinct tools to perform our analysis of synchronous vs. asynchronous settings, these tools are not representative of the full range of solutions that have been proposed to support qualitative coding. In particular, neither of these systems utilize large language models (LLMs), which have been at the forefront of recent developments in the field Dai et al. (2023); Gao et al. (2024), and which have garnered significant interest from qualitative researchers Hayes (2023); Roberts et al. (2024); Aubin Le Quéré et al. (2024).
(2) As prior work has shown Hoyle et al. (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5.3) and a user study (Sec 5.4), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, future work should expand on the human analysis of these solutions to provide a more robust view.
(3) While we look at an important variable in qualitative research settings (collaboration modality), there are several other variables that can influence the outcome of NLP-assisted solutions (e.g., choice of tool, expertise and live experience of annotators, type of data being annotated, etc.). In addition to this, we did not explore the many different consolidation strategies that are often used to bring together the perspectives of asynchronous coders. We leave the explorations of these questions for future work.
8 Ethical Considerations
To the best of our knowledge, no code of ethics was violated during the development of this project. We used publicly available tools and datasets according to their licensing agreements. For our annotation experiments, we followed IRB protocol and did not retain any personally identifiable information.
All information needed to replicate our experiments is presented in the paper. We reported all experimental settings, as well as any pre-processing steps, learning configurations, hyper-parameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results***https://github.com/blast-cu/interactive-systems .
Appendix A Interactive System Details
Operations Description
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mcinnes2017hdbscan algorithms.
Text-based Queries Experts can type any query in natural language and find instances that are close to the query in the embedding space.
Finding Similar Instances Experts have the ability to select each instance and find other examples that are close in the embedding space.
Listing Themes and Instances Experts can browse the current list of themes and their mapped instances. Instances are ranked in order of “goodness”, corresponding to the similarity in the embedding space to the theme representation. They can be listed from closest to most distant, or from most distant to closest.
Visualizing Local Explanations Experts can visualize aggregated statistics and explanations for each of the themes. To obtain these explanations, we aggregate all instances that have been identified as being associated with a theme. Explanations include wordclouds, frequent entities and their sentiments, and graphs of concept distributions.
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots vanDerMaaten2008.
(a) Exploratory Operations
Operations Description
Adding, Editing and Removing Themes Experts can create, edit, and remove themes. The only requirement for creating a new theme is to give it a unique name. Similarly, themes can be edited or removed at any point. If any instances are assigned to a theme being removed, they will be moved to the space of unassigned instances.
Adding and Removing Examples Experts can assign “good” and “bad” examples to existing themes. Good examples are instances that characterize the named theme. Bad examples are instances that could have similar wording to a good example, but that have different meaning. Experts can add examples in two ways: they can mark mapped instances as “good” or “bad”, or they can directly contribute example phrases.
Adding or Correcting Concepts We allow users to upload additional observed or predicted concepts for each textual instance. For instances and phrases added as “good” and “bad” examples, we allow users to add or edit the values of these concepts. The intuition behind this operation is to collect additional information for learning to map instances to themes.
(b) Intervention Operations
Table 5: Interactive Operations for the Pacheco et al. (2023) System

### v2

@@
-Studying the Effects of Collaboration in Interactive Theme Discovery Systems
+Effects of Collaboration on the Performance of Interactive Theme Discovery Systems
 Abstract
 NLP-assisted solutions have gained considerable traction to support qualitative data analysis.
-However, there does not exist a unified evaluation framework that can account for the many different settings in which qualitative researchers may employ them.
-In this paper, we take a first step in this direction by proposing an evaluation framework to study the way in which different tools may result in different outcomes depending on the collaboration strategy employed.
+However, no unified evaluation framework exists which can account for the many different settings in which qualitative researchers may employ them.
+In this paper, we propose an evaluation framework to study the way collaboration settings may produce different outcomes across a variety of interactive systems.
 Specifically, we study the impact of synchronous vs.
-asynchronous collaboration using two different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
+asynchronous collaboration using three different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
 11footnotetext: Equal Contribution
@@
 In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding".
-However, manual inductive coding becomes untenable as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al.
+However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al.
 (2019).
 In the last decade, several NLP solutions have been proposed to support qualitative research methods.
-These systems offer different ways to automatically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al.
+These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al.
 (2018); Fang et al.
 (2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al.
-(2023), or providing live recommendations to users as they manually code the data Dai et al.
+(2023); Chew et al.
+(2023), or providing live recommendations to users as they code the data Dai et al.
 (2023); Gao et al.
 (2024).
-To ensure the agency of the researchers, all of these systems seek to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
-In most cases, these tools are evaluated to highlight the strengths and weaknesses of the proposed system.
+To ensure the agency of the researchers, each system seeks to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
+Figure 1: In this study, we measure the quality of coded themes using different interactive systems under different coding configurations.
+Generally, tools are evaluated to highlight the strengths and weaknesses of the proposed approach.
 Examples include evaluating topic coherence with and without human input Fang et al.
-(2023) and contrasting machine-assisted code-book generation with a manual-only code-book generation Dai et al.
+(2023) and contrasting machine-assisted code-book generation with a manual code-book generation Dai et al.
 (2023).
 In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems.
-In particular, we want to study how tools for inductive coding may result in different outcomes in different collaborative settings.
-Collaborative data analysis occurs when two or more researchers pool their focus into a shared body of data to produce an agreed interpretation.
-Potential advantages of collaboration include producing more reliable results, a richer understanding of the data, and knowledge that transcends individual perspectives Flick (2014).
-While there are many ways in which collaboration can occur, in this work we look at two common but opposite qualitative research collaboration settings.
+In particular, we seek to answer the following questions: (1) Does collaboration setting measurably affect the quality of resulting code-books?
+(2) Do these findings hold across different NLP approaches?
+Collaborative Qualitative Analysis (CQA) is a process where researchers work in teams to analyze a dataset to produce a consensus interpretation.
+Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives Flick (2014).
+While there are several frameworks for conducting CQA, in this work we look at two common but contrasting collaboration settings.
 The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding).
 The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding).
-Each setting has different strengths: asynchronous coding can be a more flexible option for teams as they can work across different regions and time zones, whereas synchronous coding can combine diverse perspectives and streamline the annotation processes.
-We test these settings by choosing two NLP tools for inductive coding, and performing coding experiments in the synchronous and asynchronous contexts.
+Each setting has different strengths: asynchronous coding is more flexible as teams can work across different regions and time zones, whereas synchronous coding combines diverse perspectives and streamlines the annotation processes.
+We test these settings on three different NLP tools for interactive inductive coding by performing experiments in the synchronous and asynchronous contexts.
 We choose tools that use different underlying NLP methods for this work: (1) the Fang et al.
-(2023) system, which uses a traditional human-in-the-loop topic modeling solution, and (2) the Pacheco et al.
-(2023) system, which models emerging themes with respect to their relationship to other high-level concepts, groups similar instances based on this, and leaves theme definitions and annotations to users, which are then propagated to the full dataset.
-This way, we can have a more representative understanding of how different tools can be affected in synchronous and asynchronous collaboration.
+(2023) system, which uses a human-in-the-loop topic modeling solution, (2) the Pacheco et al.
+(2023) system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the Chew et al.
+(2023) system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset.
+By way of testing such varied approaches, we achieve a more representative understanding of how different tools perform in different collaboration settings.
 To measure outcomes in synchronous vs.
@@
 Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting.
-Finally, we also perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting.
-In summary, we make two main contributions:
-  1. 1.
-    We demonstrate how the collaboration setting in which qualitative research takes place could affect the outcome of solutions proposed in the NLP literature, by testing two NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts.
-  2. 2.
-    We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture.
-We hope that our findings will inform the future development of language technologies that are in line with the expectations and workflow of qualitative researchers, as well offer insights into how the next generation of qualitative research tools should be evaluated so that researchers are able to choose the right tool for their coding requirements.
+Finally, we perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting and make design recommendations according to our findings.
+In summary, we make two main contributions: (1) We demonstrate how the collaboration setting in which qualitative research takes place affects the outcome of solutions proposed in the NLP literature, by testing three NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts.
+(2) We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture.
+We hope that our findings will inform the development of language technologies that are in line with the qualitative researcher workflows and offer insights for producing useful qualitative research tool evaluations.
 2 Related Work
-Our work builds on previous literature on computational solutions for qualitative coding.
-The overarching goal of these systems is to partially automate the qualitative coding process by either learning user-defined themes interactively Pacheco et al.
+The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner Fang et al.
+(2023); Smith et al.
+(2018), by learning user-defined themes interactively Pacheco et al.
 (2023); Gao et al.
-(2023) or by inducing topics or clusters in an interactive, semi-supervised manner Fang et al.
-(2023); Smith et al.
+(2023), or by prompting LLMs with natural language definitions of the observed themes Chew et al.
+(2023); Dai et al.
 (2023).
 A separate but related line of work exemplified by Gao et al.
-(2024) uses language models to generate label recommendations as the user performs the coding process.
-While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to classify large portions of the dataset without requiring extensive supervision.
-Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by literature from the field of Human-Computer Interaction (HCI)  Jiang et al.
+(2024) uses LLMs to generate label recommendations as users perform the coding process.
+While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
+Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature Jiang et al.
 (2021); Feuston and Brubaker (2021); Chen et al.
 (2018).
-HCI research centers the user as the focus of study in order to identify where and how best to place NLP tools in interactive systems.
-We build on these findings by proposing a comprehensive method for evaluating a given system for qualitative coding in addition to conducting user interviews.
-To this end, we also center part of our experiment on the user to uncover differences in the interactive experience between synchronous and asynchronous environments.
-A common finding in the HCI studies surveyed is the researchers’ focus on ambiguity.
+A common finding in the HCI studies on computer-assisted qualitative research is the researchers’ focus on ambiguity.
 In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity.
@@
 With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments.
-Previous evaluation methods introduced with new systems were generally ad hoc, with experiments conducted variously in group settings  Choo et al.
+Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  Choo et al.
 (2013); Hoque and Carenini (2016); Smith et al.
@@
 3 Interactive Systems
-We evaluate two interactive theme discovery systems that use different approaches to accomplish the same task.
-This way, we ensure that our analysis can account for variation in NLP techniques and annotation interfaces.
-3.1 Fang et al. (2023)
-This system is a HitL, query-driven topic model (QDTM) (Fang et al., 2021) that allow researchers to iteratively adjust clusters using a set of refinement operations.
-Broadly, the QDTM is a semi-supervised topic modeling algorithm that accepts input queries to extract relevant concept words.
-The extracted concept words are used in a two-phase framework based off the Hierarchical Dirichlet Process (HDP) (Teh et al., 2004) to perform topic modeling.
-In the first phase, the HDP is used to infer one "parent topic" for each concept along with irrelevant topics, where each topic is defined by a set of weighted words.
-In the second phase, the model expands the parent topic into a set of subtopics using the Gibbs sampling process.
-The words for each subtopic are added to the parent topic to produce a final set of topics, which are used to assign a distribution of topics to each document.
-Given an initial topic model, users can add, remove, or swap the order of words within each topic, which alters the probability of topic assignment for documents that contain that specified word.
-Similarly, users can split or merge topics based on topic words; doing so does not guarantee that assigned documents will be reassigned accordingly due to the stochastic nature of the model.
-Users can also remove documents from a given topic, but cannot directly add documents in the same manner.
-In each iteration, users may freely select refinement operations for each topic before applying all selections at once.
-Users can undo selections prior to applying refinements and previous models are also saved if users wish to return to a previous iteration.
-The overarching goal of this system is to facilitate the discovery of an optimal topic model for a given input dataset.
-By applying a selection of refinements, users can compare the effects of different topic word allocations on the resulting model.
-Returning to previous models and applying new refinements creates a branching model, which provides users with a wide space for exploration.
-Once satisfied with a given iteration, the user can download that model with its document distribution.
-3.2 Pacheco et al. (2023)
-This system uses a two-stage framework to assist researchers in characterizing large textual collections.
-In the first stage, the system automatically proposes an initial partition of the data, such that textual instances that are thematically similar are clustered together.
-Users are then provided with a set of exploratory operations that allows them to evaluate the quality of the discovered partitions, as well as to further explore and partition the space by inspecting individual examples, finding similar instances, and using open text queries.
-Next, users are prompted to group identified patterns into general themes, and instantiate them by giving them a name.
-Once a theme is created, users are provided with a set of intervention operations to explain the themes using natural language, select good example instances, write down additional examples, and input or correct supporting concepts to characterize the theme assignments.
-The full set of operations supported by the system are outlined in App.
+We identify three categories of NLP techniques used in interactive systems for qualitative coding with large datasets: topic models, relational approaches, and LLMs.
+These techniques may be applied in a variety of ways in interactive systems, but we put special focus on their ability to help code large datasets.
+To maximize coverage across systems, we select a representative system from each category to use in our experiments.
+In this section, we briefly describe the unique aspects of each category and introduce the selected system.
+Topic Models.
+This category includes systems that use some variation of topic modeling to find emerging themes and facilitate document assignment.
+These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset.
+Early exploration incorporated visualizations to help users adjust parameters (Chuang and McFarland, 2013), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (Smith et al., 2018).
+However, topic modeling systems are limited by their lack of malleability and predictability.
+Refinement operations mostly edit topic words, which can have limited impact in the final results.
+We select the HitL query-driven topic model (QDTM) system introduced by Fang et al.
+(2023).
+The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics.
+In our experiments, users begin by iterating through each topic and naming them based on identified themes.
+Users then use a set of refinement operations to edit the topic model.
+They can merge and split topics based on topic words, add, remove, or reorder topic words, and remove documents from topics.
+The next iteration of the model is only produced when the users choose to apply refinements and the prior model is saved, allowing users to return to prior iterations to test different operations.
+Once satisfied with the state of the topic model, the user downloads the document distribution for that iteration.
+Details of our experimental settings can be found in App.
 A.
-In the second stage, the system finds a mapping between the full set of instances and the themes instantiated by the users.
-It takes the information contributed by the users in the form of examples and supporting concepts, and learns to map instances to themes using a neuro-symbolic procedure implemented with DRaiL Pacheco and Goldwasser (2021).
-The main goal of this process is to condition new theme assignments not only on the embedding distance between instances and provided examples, but also leverage the additional judgements provided by users via supporting concepts.
-For example, when analyzing a corpus about the Covid-19 vaccine, users could point out that 80% of the examples for theme Government Distrust have a clear anti-vaccine stance.
-The system then uses this information to introduce inductive bias into the mapping procedure.
-DRaiL uses weighted first-order logic rules to express decisions and dependencies between different decisions, for example $w:\mathtt{stance(anti\_vax)}\rightarrow\mathtt{theme(gov\_distrust)}$.
-For more details about this implementation, we refer the reader to the original paper.
-During this process, the systems keeps instances unassigned to a theme if there is not a good enough match.
-Then, a consecutive partitioning step is done over all unassigned instances.
-Finally, users have the option to repeat this two-stage process until convergence.
-4 Coding Protocol
-In order to study the performance of the two systems discussed in Section 3 under different collaboration settings, we design a protocol to conduct qualitative coding in the synchronous and asynchronous qualitative coding paradigms.
-We use the outcome of our coding described here to gather results, and then conduct an evaluation in the subsequent sections.
-4.1 Participants
-We recruited a group of 15 researchers in NLP and Computational Social Science, 3 female and 12 male, within the ages of 20 and 45.
-This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities.
-All participants were either well versed in qualitative data analysis, or were explicitly trained by more senior researchers to perform the task.
-4.2 Dataset
-We use a common dataset to perform qualitative coding using the Pacheco et al.
-(2023) and Fang et al.
-(2023) systems.
-This dataset was originally released by Pacheco et al.
-(2022), and consists of approximately $85,000$ Tweets about the COVID-19 vaccine.
+Relational Approaches.
+Relational approaches combine vector semantics and structured inference to model relationships between high-level concepts.
+Instead of treating themes as distributions over words (as topic models do), these frameworks define themes as distributions over generalized concepts.
+This reflects the inductive coding process, where researchers identify patterns and concepts that are then synthesized into more abstract themes.
+However, their computational complexity grows with the number of dependencies considered, which hinders their ability to quickly adapt during coding sessions.
+Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
+We select the system introduced by Pacheco et al.
+(2023), which uses a two-stage relational framework.
+In the first stage, the system automatically partitions the dataset based on semantic similarity.
+The users explore each partition to identify themes, assign "good" and "bad" example documents for each theme, and input or correct supporting concepts for each example.
+In the second stage, the system uses the provided examples and concept relations to map the remaining dataset, only leaving documents unmapped if no theme is a sufficiently good match.
+The assignment procedure follows as a structured inference approach, where dependencies between concepts and themes are explicitly modeled.
+The unmapped documents are repartitioned as in the first stage and users are prompted to review unmapped partitions again.
+The process is iterated until all documents are mapped.
+More details about the framework and our experimental configuration can be found in App.
+B
+Large Language Models.
+LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (Kojima et al., 2024).
+They have been used for theme recommendation (Gao et al., 2023), for code conflict resolution (Gao et al., 2024), and for automated document assignment (Xiao et al., 2023).
+However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering.
+Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (Chen et al., 2018).
+Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
+We select the framework introduced by Chew et al.
+(2023).
+In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code.
+The LLM is then prompted to label the data sample with the provided definitions.
+Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (Gwet, 2008).
+The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset.
+Details about our experimental settings can be found in App.
+D.
+4 Study Design
+To study the effects of different collaboration settings on the performance of the three selected systems, we design a protocol that can be used for both synchronous and asynchronous settings.
+For each system, we conduct three asynchronous experiments with one coder each and two synchronous experiments with three coders each for a total of 15 experiments.
+Evaluation metrics are calculated by comparing the resulting code-books within each experimental setting (e.g.
+the two code-books independently created by the two synchronous groups using the topic model).
+The rest of this section lays out the dataset, participant demographics and experimental protocol.
+Dataset.
+We use a common dataset for all experiments.
+It was originally released by Pacheco et al.
+(2022), and it consists of approximately $85,000$ Tweets about COVID-19 vaccines.
 All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan.
 and Oct.
-2021, and contain predictions for vaccination stance (e.g.
+2021, and contain labels for vaccination stance (e.g.
 pro-vax, anti-vax) and morality frames (e.g.
-fairness/cheating and their actor/targets.) Haidt and Graham (2007); Roy et al.
+fairness/cheating and their actor/targets.) Roy et al.
 (2021).
-4.3 Qualitative Coding Protocol
-In order to compare the consistency between different collaborative settings, we need at least two sets of synchronous/asynchronous annotations per annotation system.
-To ensure compatibility between synchronous and asynchronous outcomes, we use the same starting clusters for each system.
-The hyperparameters used in the Fang et al.
-(2023) system are the same as in their study and we did not use any input queries prior to initialization.
-We scheduled annotations for each setup separately, with details outlined in the sections below.
-4.3.1 Synchronous Coding
-Coding for the Pacheco et al.
-(2023) system was done over Zoom, where participants met online and collaboratively discussed and annotated the dataset.
-Six participants were bifurcated into two groups of three individuals, with each group communicating only with their members during the annotation process.
-Coders followed the protocol outlined in the original paper, performing two consecutive rounds of interaction, each round lasting a total of 1 hour.
-For the Fang et al.
-(2023) system, we selected a different set of six annotators and divided them into two separate groups of three individuals each.
-Annotators met with their group and collaboratively coded the data.
-Prior to the annotation session, participants were briefed about the different features of the system and how to use them.
-The annotators went over the pre-assigned topics proposed by the system, and modified them to make them more accurate and cohesive.
-Both groups of annotators reached saturation after multiple passes over the data, in sessions lasting around 2.5 hours.
-4.3.2 Asynchronous Coding
-The asynchronous coding processed followed the same outline as the synchronous coding for each system, except that it was conducted by three different annotators, and it was done individually without any collaboration.
-For each system, the annotators first met in-person and started annotations separately.
-The purpose of the initial co-located coding session was to troubleshoot any problems and solve any doubts that the participants may have.
-The participants then completed the annotation completely asynchronously and remotely.
-Pacheco et al. (2023) Fang et al. (2023)
-Jaccard Sim. Centroid Sim. Group Avg. Sim. Jaccard Sim. Centroid Sim. Group Avg. Sim.
-Sync $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$
-Asyn $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$
+Participants.
+We recruited a group of 20 researchers in NLP and Computational Social Science, 6 female and 14 male, between the ages of 20 and 45.
+This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities.
+This group covers the range of researchers likely to use interactive coding systems.
+All participants were either well-versed in qualitative data analysis, or were explicitly trained by senior researchers to perform the task.
+Coding Protocol.
+At the start of each experiment, participants were provided with a demonstration of all the operations in their respective systems.
+Every system starts with an initial partition of the data, so participants were instructed to read the first 25 samples in each partition, and manually create/name any themes they identified before freely exploring the rest of the dataset and start performing operations to find more themes.
+In the topic model experiments, we suggested that participants merge and split topics based on their identified themes before making fine-grained refinements.
+They were then asked to refine the topic model based on their identified themes such that every topic corresponds to a unique theme.
+They kept re-running the model and making refinements until they were satisfied with the results, or until they failed to effect any meaningful changes.
+In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them.
+Following Pacheco et al.
+(2023), the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame).
+Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset.
+Unmapped examples were repartitioned and returned to the participants for a second (and last) round of coding.
+In the experiments for the LLM-based system, participants produced natural-language definitions for each identified theme and selected a set of good examples for them.
+We then prompted the LLM with different task-prompt templates to find the best prompt for each set of participant-generated definitions, which was then used to code the rest of the dataset.
+Details of the templates, as well as the human-model agreement for the best template can be found in Appendix D.
+Topic Model Relational LLM-Based
+Jaccard Centroid Group Avg. Jaccard Centroid Group Avg. Jaccard Centroid Group Avg.
+Sync $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$ $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $0.14(0.08)$ $0.98(0.03)$ $0.44(0.03)$
+Async $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$ $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $\bf 0.17(0.11)$ $0.98(0.02)$ $\bf 0.45(0.03)$
 Table 1: Avg.
@@
 5 Evaluation
-In this section, we use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings.
-We first define and present several metrics for theme induction quality based on overlap, similarity/distance, association and verification that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al.
+We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings.
+Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al.
 (2021); Pacheco et al.
 (2023).
-These metrics are calculated using the sets of documents assigned to each theme by their respective system.
-While the Pacheco et al.
-(2023) system explicitly assigns a single theme for each document, the Fang et al.
-(2023) system provides a distribution of theme weights for each document.
-To ensure that results from both systems are comparable, we assign each document to the theme with the highest weight.
-Finally, we summarize annotator experiences to provide a perspective on qualitative differences between the two settings.
-Combining dimensions in our evaluation methodology, we demonstrate that there are significant differences in consistency, cohesiveness, correctness, and user experiences between synchronous and asynchronous experiments.
-5.1 Consistency
-Coding individually always runs the risk that coders overgeneralize or miss key themes, resulting in “unconstrained or unsystematic” analysis (Cornish et al., 2014).
-We propose consistency as one dimension for measuring this systematicity.
-Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts.
-There are two obstacles to measuring annotation consistency when using semi-automated annotations.
-First, themes with similar names may not share the same annotated documents and themes with different names may share documents that are semantically similar, which we address by measuring the consistency of documents between given pairs of themes.
-Following the definition of consistency in Section 4 of Pacheco et al.
-(2023), we calculate the Jaccard similarity given as:
-$J(A,B)=\frac{|A\cap B|}{|A|+|B|-|A\cap B|}$ (1)
-where $A$ and $B$ are sets of documents for two themes with different coders.
-We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
-Second, since the majority of documents are assigned a theme by the system, the same themes may contain a different assortment of documents but remain relatively semantically similar.
-We therefore also measure the semantic similarity between themes using their SBERT embeddings Reimers and Gurevych (2019).
+Figure 2: Two sets of annotators use a particular HiTL system to find themes.
+Since the same theme can be named differently by different annotators, we find the best match.
+In this example, the annotator 1’s theme "VaxIsBad" has been matched with annotator 2’s theme "antivax".
+After aligning, we calculate the similarity between these two themes using methods like Jaccard Similarity or Centroid Distance.
+Consistency.
+Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (Cornish et al., 2014).
+We propose consistency as one dimension for measuring systematicity.
+Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts (See Fig.
+2).
+Once the dataset has been annotated using an interactive system, there are two obstacles to measuring consistency when using semi-automated annotations.
+First, themes with similar names may not share the same annotated documents.
+Similarly, themes with different names may share the same documents.
+We measure consistency using the overlap of documents between a given pairs of themes.
+To do this, we calculate the Jaccard similarity across the documents for two themes with different coders.
+We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best aligned theme”.
+In Fig.
+2, this is shown in the middle where the colored lines indicate the best match obtained by maximum Jaccard overlap.
+After alignment, we calculate the average of all matched pairs for comparison across settings.
+We want to account for the possibility of themes containing a different assortment of documents that otherwise remain relatively semantically similar.
+To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them Reimers and Gurevych (2019).
 We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme.
 (2) Group avg.
-similarity, where the similarity between two themes is the average of pair-wise similarity between all documents in the two themes.
-As in the Jaccard case, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
-In all cases, we use the cosine similarity.
+similarity, where the similarity between two themes is the average of pair-wise cosine similarity between all documents in the two themes.
+As in the Jaccard metric, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
 Table 1 shows the average maximum Jaccard and embedding similarities for each experiment.
-Comparing the results on the synchronous/asynchronous dimension, we find that synchronous groups produce themes with higher average consistency across the board.
-The centroid similarities and group average similarities in particular are statistically significant, supporting our hypothesis that different settings produce noticeably different coding results.
-These findings also demonstrate that the theoretical advantage for collaborative coding holds empirically true when using HitL systems as well: coders are more likely to miss themes or overgeneralize when working independently.
-Comparing between consistency metrics, we find that centroid similarity is at or near statistical significance for all experiments and is much higher than the other metrics across the board.
-This finding is also in line with our expectations detailed above: since HitL systems automate a large portion of document assignments, small differences in user interactions can result in larger differences in document mappings.
-However, Jaccard similarity is still important to consider, especially for datasets comprised of documents that are very semantically similar.
-5.2 Cohesiveness and Distinctiveness
-Pacheco et al. (2023) Fang et al. (2023)
-Intra-Theme Inter-Theme Intra-Theme Inter-Theme
-Sync $\bf 0.51(0.08)^{*}$ $0.41(0.04)^{*}$ $0.52(0.10)$ $0.40(0.04)$
-Async $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.52(0.10)$ $0.40(0.04)$
+We find that synchronous groups produce themes with higher average consistency for the topic modeling and relational systems.
+For the LLM-based system, we did not observe statistical significant differences across collaboration settings.
+Interestingly, the LLM-based system offered the least amount of intervention/refinement operations for users during the coding process.
+Whereas the other systems allowed for operations such as splitting and merging topics (topic modeling), and defining relations between concepts (relational approach), the LLM system only allowed users to alter the definition of the theme provided in the prompt.
+We hypothesize that the more interventions users can make, the more advantage they can take of the deliberation afforded by synchronous coding.
+Topic Model Relational LLM-based
+Intra-Theme Inter-Theme Intra-Theme Inter-Theme Intra-Theme Inter-Theme
+All Sync $0.52(0.10)$ $0.40(0.04)$ $\bf 0.51(0.08)^{*}$ $0.42(0.05)^{*}$ $\bf 0.44(0.06)$ $0.40(0.04)$
+Async $0.52(0.10)$ $0.40(0.04)$ $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.43(0.05)$ $\bf 0.39(0.04)$
+Top 25% Sync $0.56(0.11)$ $0.39(0.05)$ $\bf 0.70(0.09)^{*}$ $0.52(0.07)^{*}$ $0.63(0.07)$ $0.55(0.05)$
+Async $0.56(0.11)$ $0.39(0.05)$ $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.63(0.05)$ $\bf 0.54(0.05)$
 Table 2: Group Avg.
-Similarity within Annotator Groups (All data points).
-Themes are considered to be more cohesive if intra-theme similarity is high, and more distinctive if inter-theme similarity is low.
+Similarity within Annotator Groups.
+Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low.
 Statistically significant using a two-sample unpaired t-test with $p<0.05$.
-A second dimension for determining the "systematicity, clarity, and transparency" of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book (Cornish et al., 2014).
-We propose two metrics, cohesiveness and distinctiveness, to represent these similarities and differences respectively.
-Whereas consistency measures the systematicity and clarity of themes across annotators, cohesiveness and distinctiveness measure these qualities for themes by the same annotator.
-The distinctiveness of a theme is determined by comparing it to other themes created by the same annotator.
-On the other hand, the cohesiveness of a single theme is determined by its semantic homogeneity.
+Figure 3: Once an annotator has identified themes and they have been propagated the full dataset, we calculate intra-theme similarity by measuring the avg.
+of the pairwise distances between each document within a theme (left).
+We calculate inter-theme similarity by measuring the avg.
+of pairwise distances between each document in a theme and documents assigned to all other themes (right)
+Cohesiveness and Distinctiveness.
+Another dimension for determining the systematicity and clarity of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book.
+We propose two metrics to measure this: cohesiveness and distinctiveness.
 A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity).
@@
 If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
-To evaluate these metrics for a single annotator (or annotator group), we calculate both the intra-theme similarity and the inter-theme similarity for theme in the code-book.
+Figure 3 shows how to evaluate these metrics for a single annotator (or annotator group).
+We calculate both the intra-theme similarity and the inter-theme similarity for all the themes in the code-book.
 Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme.
-Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes by the same annotator.
-A confounding factor in these measures is that both systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme.
-In order to more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme.
-For the Pacheco et al.
-(2023) system, this top quartile is selected using the distance from the centroid.
-For the Fang et al.
-(2023) system, the distribution of theme weights better reflect the assignments provided by the topic model and are used instead.
-Table 2 shows results for the whole dataset and Table 3 show the same metrics for the subset of the documents closest to each theme.
+Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes.
+A confounding factor in these measures is that all systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme.
+To more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme.
+For the relational and LLM-based systems, this top quartile is selected using the distance from the centroid.
+For the interactive topic model, we use the weights assigned by the model.
+Table 2 shows results for both the whole dataset as well as the subset of the documents closest to each theme.
 Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board.
-For the Pacheco et al.
-(2023) experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting.
-Our findings hint that themes may be more cohesive but not more distinctive in the synchronous setting.
-Pacheco et al. (2023) Fang et al. (2023)
-Intra-Theme Inter-Theme Intra-Theme Inter-theme
-Sync $\bf 0.70(0.08)^{*}$ $0.52(0.07)^{*}$ $0.56(0.11)$ $0.39(0.05)$
-Async $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.56(0.10)$ $0.39(0.05)$
-Table 3: Group Avg.
-Similarity within Annotator Groups (25% most related).
-Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low.
-Statistically significant using a two-sample unpaired t-test with $p<0.05$.
-The results in the Fang et al.
-(2023) experiments are nearly identical across collaboration modalities.
-Narrowing down to the top 25% of closest documents increases the difference between the two metrics, lining up with the intuitive notion that doing so increases both cohesiveness and distinctiveness.
-When comparing the two systems, we find that the themes coded using the Pacheco et al.
-(2023) system are significantly more cohesive when looking at the quartile subset, which matches our findings in Section 5.3.
-5.3 Correctness
-While the previous evaluations look at the code-book as a hole, we want to also evaluate how correct our document to theme assignments are.
-In order to verify this, we conducted a post-hoc analysis of how accurate theme assignments are on a held-out set of randomly sampled tweets, compared across synchronous and asynchronous settings.
-The data was sampled after the annotation process was over.
-This meant that for both systems, annotators had deemed the theme discovery and refinement to have reached saturation.
-Following Section 5.2, we perform one analysis of correctness that considers the set of all tweets, as well as a second analysis considers only the top $25\%$ of tweets most closely related to each theme.
-For each analysis, we sampled $200$ tweets randomly and had two of the authors of the paper manually annotate whether the theme assigned to a tweet by the system makes sense or not.
-We did this for each system and collaboration setting (synchronous vs.
-asynchronous), with the exception of the Pacheco et al.
-(2023) synchronous setting, for which we report the results presented in the original paper.
-We made sure to follow the same sampling procedure so that results were comparable.
-Results are presented in Tab.
-4.
-We observe that for both systems, the accuracy yielded from synchronous annotations is higher than their asynchronous counter parts.
-We also see that as per our expectations, the top $25^{th}$ percentile has a higher accuracy than group average of all tweets.
-However, it is interesting to observe the comparatively poor accuracy of the topic modeling system Fang et al.
-(2023) as compared to the theme discovery system Pacheco et al.
-(2023).
-We attribute this difference to the fact that topic modeling systems do not "learn" to classify themes, but rather learn the probability that a given word belongs to a particular topic.
-During evaluation, we found that the main problem arises when salient words are used in atypical contexts, or when trying to bifurcate terms that are used in opposing themes.
-As an example, take the following headline regarding the COVID-19 vaccine mandate.
-Indiana University students must comply with the school’s COVID-19 vaccine mandate for the fall semester after a federal judge refused on Monday to block the school’s inoculation requirement, one of the first rulings of its kind during the pandemic
-The topic modeling system gave the highest wright to the topic Anger over mask mandates.
-But really, this is regarding vaccines.
-The term mandate biases the model to rank a topic about masks (the context that typically appears around the word "mandate") higher than a topic related to the COVID-19 vaccine, which this headline is about.
-Our findings regarding the correctness can be summarized by acknowledging that synchronous teams of annotators achieve significantly higher assignment accuracy than asynchronous teams.
-We also find that topic modeling systems achieve un-competitive assignment accuracy versus theme discovery systems.
-Lastly, qualitative researchers may want to choose theme discovery tools if they want higher accuracy in document to theme assignments – but could opt for topic modeling systems if they want to put their focus on word usage patterns.
-Pacheco et al. (2023) Fang et al. (2023)
-All Top 25% All Top 25%
-Sync ${}^{*}68\%$ ${}^{*}86\%$ $23\%$ $42\%$
-Async $50\%$ $64\%$ $17\%$ $35\%$
-Table 4: Correctness: Accuracy of Theme Assignment for Held-out Dataset. Results taken from Pacheco et al. (2023).
-5.4 User Study
-We carried out semi-structured interviews with all the participants in the study after they were done with their coding task.
+For the relational system experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting.
+Our findings suggest that themes may be more cohesive but not more distinctive in the synchronous setting.
+Results for the interactive topic model and LLM-based systems are nearly identical across collaboration modalities.
+Given that the operations to characterize themes offered by these two systems are substantially limited (i.e., adding and removing words for the topic model, and refining theme definitions for the LLM system), these results are not surprising.
+Again, we see that the wider the range of operations afforded to characterize a theme, the more advantage can be taken from synchronous deliberation.
+Correctness.
+Interactive systems allow users to automate large portions of the coding process at the risk of producing inaccurate theme assignments.
+To estimate how correct the outputs of each system are, we conduct a post-hoc analysis by manually checking a randomly selected sample of $1,200$ document-theme pairs ($200$ per experimental setting).
+To ensure that our sample is representative of the overall dataset for each experimental setting, we split the data into quartiles based on document relatedness to each theme and select a uniform sample of themes and relatedness scores.
+As in the previous section, relatedness is calculated using the theme weight distribution for the interactive topic model and distance from the centroid for the the other two systems.
+To assess reliability, each assignment is evaluated by two annotators, with a third weighing in for tie breaks.
+We also find that human evaluators have moderate-to-high agreement when assessing system outputs (with an overall Krippendorff’s $\alpha$ of $0.632$), suggesting that we can trust these estimations.
+Figure 4 shows the correctness results for each quartile sample per experiment.
+First, we observe that the relational system is not only the most accurate, but it shows negligible correctness differences in synchronous vs.
+asynchronous configurations.
+This an encouraging result, given that this system took the most advantage of synchronous deliberation based on our other metrics of quality.
+The other two systems showed marked differences, with the topic modeling approach producing more accurate assignments in asynchronous operation, and the LLM producing more accurate assignment in a synchronous paradigm.
+This observation leads us to believe that interactive systems that introduce less and weaker inductive biases from interaction are more sensitive to choices of configuration – and researchers should be aware of these possibilities when designing experiments.
+Second, we find that the correctness of different approaches tapers differently, based on the distance of the example from the theme.
+In other words, as the data samples resemble less like the human labeled ones, some models handle it more robustly than others.
+In our observation, relational approaches and LLM methods outperform topic models when it comes to assigning themes to distant examples.
+Figure 4: Correctness w.r.t. distance from theme.
+6 User Study and Recommendations
+We carried out semi-structured interviews with the participants in the study.
 We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities.
+Interview questions can be found in App.
+C.
 We summarize our main findings below.
-Synchronous annotators found working in teams to ease the annotation process.
-Both teams of annotators working synchronously using the Fang et al.
-(2023) described their perceived advantage of working in teams – broader perspectives by inviting view points from different annotators helped them contextualize diverse tweets easily.
-The annotators also valued the ability to collaboratively resolve contested tweets through discussions.
-Particularly, it was easy to "break ties" quickly through conversation when it came to mapping which words belonged to a given a topic.
-The annotators were also quicker to reach saturation, and they expressed that they thought this was possible because consolidation was possible on-the-fly.
-One of the annotators noted that a pre-requisite for successful synchronous coding stemmed from the rapport their annotation team shared, as they are all members of the same lab and were able to convey their viewpoints without hesitation.
-Asynchronous annotators were more likely to notice inadequacies of the annotation tool they used.
-While most of the highlights from the synchronous annotators focused on the process of annotation, the asynchronous annotators directed their feedback towards the way the tools functioned.
+Participants found working in teams eases the coding process.
+Teams of annotators working synchronously described their perceived advantage of working in teams – broader perspectives helped them contextualize diverse tweets easily.
+The annotators also valued the ability to collaboratively resolve disagreements through discussions.
+They cited it being easy to "break ties" quickly through conversation.
+In addition to easing the user experience, we also note that the consistency and cohesiveness of annotations is higher in synchronous settings.
+Based on this evidence, future work should consider designing systems that explicitly support synchronous deliberation.
+Asynchronous coders were more likely to notice inadequacies of the tool they used.
+Asynchronous annotators directed most of their feedback towards the way the tools functioned.
 Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts.
-The synchronous annotators also met with similar issues, but were quick to solve it collaboratively.
-Even though all annotators (in both settings and systems) started with the same initial clusters of tweets, the synchronous annotators were able to find the operation they need performed quicker through in-team collaboration.
-Annotators noted the lack of control and agency that the topic modeling system had over the theme discovery system.
+Based on these findings, tool designers should focus on improving the asynchronous coding experience.
+Annotators noted a lack of control and agency in the topic modeling system.
 Annotators expressed loss of control and agency when using the topic modeling system Fang et al.
-(2023) over the theme discovery system Pacheco et al.
-(2023).
+(2023) over other systems.
 One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups".
+Based on these findings, tool designers should consider abstracting users away from the intricacies of the underlying NLP method.
 Annotators felt that starting topics often negatively impacted their annotations.
-Since topic models start by a sampled cluster of data, both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic.
+Both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic in the topic modeling system.
+In contrast to the clustering approach used by the other two systems, which appeared to yield cleaner partitions.
 One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al.
 (2023) model groups together despite the stark difference in message between the two.".
-Since the topics contain a large number of documents, it is only feasible for annotators to include / exclude a very small subset as being part of the topic.
 Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended.
-However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of clustering already done - as opposed to the Pacheco et al.
-(2023) system where annotations had to be done from scratch, posing to be a small hurdle in the beginning.
+However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of theme induction done already - as opposed to the other two systems, where clusters did not have any representative words attached to them.
+Based on these findings, researchers should investigate ways to effect more change during interactions for topic modeling systems, as well as consider generating some initial explanations to characterize clusters for alternative approaches.
+Using LLMs to conduct thematic coding at scale is expensive and less reliable.
+Despite the recent advances in the reasoning capabilities of LLMs, we observe that their performance at scale is not effective enough to be used for the purpose of propagating annotations.
+Additionally, running LLMs for the $5$ annotation configurations came with significant time and resource costs.
+This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners, and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
+Future work developing interactive systems should take into account the amount of control the system offers to the user.
+Quantitatively, we find that the topic modeling system produced results that were less correct across the board.
+This matches with our user study where users were unable to disentangle themes that had a high degree of lexical similarity.
+Despite being able to refine the topic model such that the highest weighted documents matched their defined themes, users had little control over whether or not the rest of the documents in the dataset were faithfully assigned.
 7 Conclusion and Future Work
-In this paper, we examined NLP-assisted qualitative research tools in different collaboration modalities.
-We conducted annotations on a vast dataset using two popular interactive coding tools, Pacheco et al.
-(2023) and Fang et al.
-(2023).
-We then designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration.
+In this paper, we examined three categories of NLP-assisted qualitative research tools in different collaboration modalities, and conducted inductive and deductive coding on a large dataset of English texts.
+We designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration.
 We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output.
-This is particularly true for theme discovery systems like Pacheco et al.
-(2023), where users have a wider bandwidth of interaction that can benefit from group consensus.
-We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and highly accurate document-level annotations.
-While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy in supporting qualitative coding.
-We believe that the evaluation framework that we proposed can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated.
+This is particularly true for systems like Pacheco et al.
+(2023), where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding.
+We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments.
+Finally, we show that while LLM-based solutions show promise, they pose significant challenges when it comes to coding at scale.
+While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy for qualitative coding.
+We believe that our proposed evaluation framework can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated.
 Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.
 8 Limitations
-The study presented in this paper has three main limitations.
-(1) While we selected two distinct tools to perform our analysis of synchronous vs.
-asynchronous settings, these tools are not representative of the full range of solutions that have been proposed to support qualitative coding.
-In particular, neither of these systems utilize large language models (LLMs), which have been at the forefront of recent developments in the field Dai et al.
-(2023); Gao et al.
-(2024), and which have garnered significant interest from qualitative researchers Hayes (2023); Roberts et al.
-(2024); Aubin Le Quéré et al.
-(2024).
+The study presented in this paper has four main limitations.
+(1) While we selected three distinct, representative tools to perform our analysis of synchronous vs.
+asynchronous settings, the list is of course non-exhaustive.
+A larger study incorporating more tools could yield additional insights.
 (2) As prior work has shown Hoyle et al.
 (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling.
-Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5.3) and a user study (Sec 5.4), both of these evaluations were limited in scope.
+Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope.
 This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation.
-While we are confident in the findings of our study, future work should expand on the human analysis of these solutions to provide a more robust view.
+While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
+(3) We used a single dataset in our experiments, which is composed of short English texts.
+While we believe that this was sufficient to arrive at strong conclusions across systems and collaboration modalities, datasets with different characteristics (e.g., longer documents, non-English language) could pose additional challenges.
 (4) While we look at an important variable in qualitative research settings (collaboration modality), there are several other variables that can influence the outcome of NLP-assisted solutions (e.g., choice of tool, expertise and live experience of annotators, type of data being annotated, etc.).
@@
 Due to space constraints, some of this information was relegated to the Appendix.
-In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results***https://github.com/blast-cu/interactive-systems .
-Appendix A Interactive System Details
+In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results..
+Appendix A Topic Model Experimental Settings
+To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al.
+(2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$.
+The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability.
+The same initial topic model is provided for all experiments.
+Appendix B Experimental Settings and Interactive System Details
+During the interactive coding process, researchers are provided initial clusters to identify themes.
+We use $K=10$ means clustering to generate initial partitions and the same partitions are provided for all experiments.
+Users select positive and negative examples for each theme, which are used to create a distributional representations to calculate semantic similarity with unlabeled documents.
+Simultaneously, supporting concepts are defined using logical rules which researchers can use to label example documents.
+The relationship between themes and concepts are then used as a structured inference task to predict label assignments.
+By using multiple sources of information, the model can much more efficiently extrapolate from a small set of manually labeled data.
+B.1 Relational Approach Operational Details
 Operations Description
-Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mcinnes2017hdbscan algorithms.
+Finding Partitions Experts can find partitions in the space of unassigned instances.
+We currently support the K-means Jin and Han (2010) and Hierarchical Density-Based Clustering Mendelsohn et al.
+(2021) algorithms.
 Text-based Queries Experts can type any query in natural language and find instances that are close to the query in the embedding space.
@@
 To do this, we aggregate all instances in the database.
-Explanations include theme distribution, coverage statistics, and t-sne plots vanDerMaaten2008.
+Explanations include theme distribution, coverage statistics, and t-sne plots van der Maaten and Hinton (2008).
 (a) Exploratory Operations
@@
 Table 3: Interactive Operations for the Pacheco et al. (2023) System
+Appendix C Semi-Structured Interview
+C.1 Interviewing
+We usually administered interviews after annotation sessions.
+In some cases, due to time constraints from the annotation session taking longer than expected, we were unable to pose these questions to the annotators.
+Asynchronous annotators were asked questions individually about their experience, whereas synchronous annotator groups were usually asked questions with their fellow annotators.
+C.2 Script
+  1. 1.
+    Have you worked on annotation projects before? Did these annotation projects use qualitative coding strategies (ex: grounded theory)? How experienced are you as an annotator?
+  2. 2.
+    How was your experience on the COVID-19 vaccine annotation session we conducted on Sunday? Particularly, we are interested in your thoughts and feelings over the session.
+  3. 3.
+    You annotated in a group, working together as a team.
+Did you find this setup to be beneficial?
+What were some of the limitations you faced, both individually and and as a group, when working synchronously?
+  4. 4.
+    On a similar line, what would you consider to be the pros and cons if you were to annotate alone?
+The last questions would be flipped based on if we are posing it to synchronous or asynchronous annotators.
+Appendix D LLM-based Experimental Configuration and Prompt Details
+For this study, we use the same starting partitions as used in the relational approach experiment in App.
+B.
+The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
+code-book Gwet’s $AC_{1}$ Unlabeled Docs
+Sync 1 $0.42$ $5,548(6.5\%)$
+Sync 2 $0.49$ $9,766(11.4\%)$
+Async 1 $0.61$ $611(0.7\%)$
+Async 2 $0.62$ $2,506(2.9\%)$
+Async 3 $0.46$ $13,816(16.1\%)$
+Table 4: Results for the selected prompt for each coding session using the LLM-based system.
+Gwet’s $AC_{1}$ is used to select the best prompt for running the full dataset.
+The number of unlabeled documents represent documents where the LLM produced a label not created by human annotators after running the full dataset (percentage of the dataset unlabeled).
+Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work.
+We used the Llama 3.2 3B-Instruct (Grattafiori et al., 2024) model for all generation tasks:
+D.1 LLM Hyperparameters
+Batch size: 32
+Model: Llama 3.2 3B-Instruct
+GPU: A100 40GB VRAM
+Average Compute Time: 24hrs per job
+Number of jobs: 5
+D.2 LLM Prompts
+To code this tweet, do the following:
+- First, read the codebook and the tweet.
+- Next, decide which code is most applicable and explain your reasoning for the coding decision.
+- Finally, generate json with your code and your reason for the coding decision. The respone MUST be formatted as JSON.
+Codes:
+<codes>
+Codebook:
+<codebook>
+Tweet:
+<tweet>
+JSON Output:
+"code" : "",
+"reason" : ""
+To code this tweet, do the following:
+First, read the codebook and the tweet.
+Next, decide which code is most applicable based on the tweet’s content and explain your reasoning for the coding decision.
+Finally, generate a JSON object with the selected code and provide a brief explanation for your coding decision.
+The response MUST be formatted as JSON.
+Codebook: Themes: <"theme": "definition">
+Tweet: < "text": "<text>" >
+JSON Output: < "code": "", "reason": "" >
+To generate code for this tweet, provide a step-by-step explanation of how to approach the task.
+First, analyze the tweet’s content and identify key concepts, such as the type of object or class being described, any specific behaviors or requirements, and relevant keywords.
+Next, evaluate the codebook options and determine which one is most applicable.
+Explain your reasoning for your decision, including any similarities between the tweet and the code definitions, or any specific requirements mentioned in the tweet that align with a particular code.
+Finally, generate a JSON object with the selected code and provide additional context, including:
+A clear explanation of how you arrived at your chosen code
+Any relevant notes or comments about the code’s functionality and requirements
+A brief comparison to other codes in the book, if applicable
+The response MUST be formatted as JSON.
+Codebook: <codebook>
+Tweet: <tweet>
+JSON Output: < "code": "", "reasoning": "", "context": "" >
+To analyze this tweet and select a relevant theme, follow these steps:
+First, read the tweet and identify key concepts, such as emotions, objects, or ideas mentioned in the text.
+Next, evaluate the theme options and determine which one is most applicable.
+Explain your reasoning for your decision, including any connections you see between the tweet’s content and the theme definitions.
+Then, generate a JSON object with the selected theme and provide additional insight into your analysis. Include:
+A clear explanation of how you arrived at your chosen theme
+Any specific characteristics or keywords from the tweet that support your decision
+A brief comparison to other themes, if applicable
+The response MUST be formatted as JSON.
+Themes: <Codebook>
+Tweet: <tweet>
+JSON Output: <"theme": "", "insight": "">

### v3

@@
 1 Introduction
-Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968).
+Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (braun-2006) and Grounded Theory (glaser1968discovery).
 In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding".
-However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al.
-(2019).
+However, manual inductive coding becomes infeasible as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process doi:10.1146/annurev-polisci-090216-023229; IJoC10675.
 In the last decade, several NLP solutions have been proposed to support qualitative research methods.
-These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al.
-(2018); Fang et al.
-(2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al.
-(2023); Chew et al.
-(2023), or providing live recommendations to users as they code the data Dai et al.
-(2023); Gao et al.
-(2024).
+These systems offer different ways to systematically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models smith-2018-loop; fang-2023-user, grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset pacheco-2023-interactive; chew2023llmassistedcontentanalysisusing, or providing live recommendations to users as they code the data dai-etal-2023-llm; gao-2024-collabcoder.
 To ensure the agency of the researchers, each system seeks to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
@@
 Generally, tools are evaluated to highlight the strengths and weaknesses of the proposed approach.
-Examples include evaluating topic coherence with and without human input Fang et al.
-(2023) and contrasting machine-assisted code-book generation with a manual code-book generation Dai et al.
-(2023).
+Examples include evaluating topic coherence with and without human input fang-2023-user and contrasting machine-assisted code-book generation with a manual code-book generation dai-etal-2023-llm.
 In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems.
@@
 Collaborative Qualitative Analysis (CQA) is a process where researchers work in teams to analyze a dataset to produce a consensus interpretation.
-Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives Flick (2014).
+Doing so can produce more reliable results, a richer understanding of the data, and insight that combines multiple perspectives cite-key.
 While there are several frameworks for conducting CQA, in this work we look at two common but contrasting collaboration settings.
@@
 We test these settings on three different NLP tools for interactive inductive coding by performing experiments in the synchronous and asynchronous contexts.
-We choose tools that use different underlying NLP methods for this work: (1) the Fang et al.
-(2023) system, which uses a human-in-the-loop topic modeling solution, (2) the Pacheco et al.
-(2023) system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the Chew et al.
-(2023) system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset.
+We choose tools that use different underlying NLP methods for this work: (1) the fang-2023-user system, which uses a human-in-the-loop topic modeling solution, (2) the pacheco-2023-interactive system, which models emerging themes with respect to their relationship to other high-level concepts which can be defined by users, and (3) the chew2023llmassistedcontentanalysisusing system, which leverages large language models (LLMs) to judge the adequacy of code definitions, as well as to broadcast them to a large dataset.
 By way of testing such varied approaches, we achieve a more representative understanding of how different tools perform in different collaboration settings.
@@
 2 Related Work
-The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner Fang et al.
-(2023); Smith et al.
-(2018), by learning user-defined themes interactively Pacheco et al.
-(2023); Gao et al.
-(2023), or by prompting LLMs with natural language definitions of the observed themes Chew et al.
-(2023); Dai et al.
-(2023).
-A separate but related line of work exemplified by Gao et al.
-(2024) uses LLMs to generate label recommendations as users perform the coding process.
+The overarching goal of the systems we investigate is to partially automate the qualitative coding process either by inducing topics in an interactive, semi-supervised manner fang-2023-user; smith-2018-loop, by learning user-defined themes interactively pacheco-2023-interactive; gao-2023-coaicoder, or by prompting LLMs with natural language definitions of the observed themes chew2023llmassistedcontentanalysisusing; dai-etal-2023-llm.
+A separate but related line of work exemplified by gao-2024-collabcoder uses LLMs to generate label recommendations as users perform the coding process.
 While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to annotate large portions of the dataset without extensive supervision.
-Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature Jiang et al.
-(2021); Feuston and Brubaker (2021); Chen et al.
-(2018).
+Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by the Human-Computer Interaction (HCI) literature jiang-2021-serendipity; feuston2021tools; chen-2018-ambiguity.
 A common finding in the HCI studies on computer-assisted qualitative research is the researchers’ focus on ambiguity.
@@
 However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas.
-Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size.
+Solutions have been proposed to either visualize codes (drouhard-aeonium) or rank document disagreement (zade-2018-disagree) regardless of dataset size.
 Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations.
@@
 With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments.
-Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  Choo et al.
-(2013); Hoque and Carenini (2016); Smith et al.
-(2018), on individual participants  Rietz et al.
-(2020), and through platforms such as MTurk  Zade et al.
-(2018).
+Previous evaluation methods introduced with new systems have generally been ad hoc, with experiments conducted variously in group settings  choo-utopian; hoque-convisit; smith-2018-loop, on individual participants  rietz-cody, and through platforms such as MTurk  zade-2018-disagree.
 Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
@@
 These systems benefit from the relative speed of the topic model, which allow users to quickly visualize and explore the dataset.
-Early exploration incorporated visualizations to help users adjust parameters (Chuang and McFarland, 2013), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (Smith et al., 2018).
+Early exploration incorporated visualizations to help users adjust parameters (Chuang2013DocumentEW), while later works implemented refinement operations that allow users to directly edit topic words and remove documents (smith-2018-loop).
 However, topic modeling systems are limited by their lack of malleability and predictability.
 Refinement operations mostly edit topic words, which can have limited impact in the final results.
-We select the HitL query-driven topic model (QDTM) system introduced by Fang et al.
-(2023).
+We select the HitL query-driven topic model (QDTM) system introduced by fang-2023-user.
 The topic model is initialized by providing input queries (i.e., words that represent concepts of interest for the user) which the model uses to generate the initial topics.
@@
 Further, they rely on users to define informative concepts, making them less suited for inexperienced researchers.
-We select the system introduced by Pacheco et al.
-(2023), which uses a two-stage relational framework.
+We select the system introduced by pacheco-2023-interactive, which uses a two-stage relational framework.
 In the first stage, the system automatically partitions the dataset based on semantic similarity.
@@
 Large Language Models.
-LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (Kojima et al., 2024).
-They have been used for theme recommendation (Gao et al., 2023), for code conflict resolution (Gao et al., 2024), and for automated document assignment (Xiao et al., 2023).
+LLMs are ideal candidates for interactive systems, especially for tasks such as qualitative coding where the model can be prompted to produce themes or explanations without ad-hoc training (llms0shotreasoners).
+They have been used for theme recommendation (gao-2023-coaicoder), for code conflict resolution (gao-2024-collabcoder), and for automated document assignment (deductivecodingwithLLM).
 However, the flexibility of LLM outputs also leads to hallucinations, which are only partially addressed by prompt engineering.
-Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (Chen et al., 2018).
+Models further suffer from biases in training which are difficult to identify and impact their ability to produce quality labels or recommendations (chen-2018-ambiguity).
 Additionally, their massive size is prohibitive when working with large datasets due to the high cost of inference.
-We select the framework introduced by Chew et al.
-(2023).
+We select the framework introduced by chew2023llmassistedcontentanalysisusing.
 In their protocol, the human coder first manually codes a representative subset of the data and drafts definitions for each code.
 The LLM is then prompted to label the data sample with the provided definitions.
-Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (Gwet, 2008).
+Agreement is calculated between human and model annotations using GWET’s $AC_{1}$ (gwet2008computing).
 The prompt is then tweaked iteratively to achieve a satisfactory level of agreement, and the best-performing version is used to prompt the model to code the rest of the dataset.
@@
 We use a common dataset for all experiments.
-It was originally released by Pacheco et al.
-(2022), and it consists of approximately $85,000$ Tweets about COVID-19 vaccines.
+It was originally released by pacheco-etal-2022-holistic, and it consists of approximately $85,000$ Tweets about COVID-19 vaccines.
 All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan.
@@
 pro-vax, anti-vax) and morality frames (e.g.
-fairness/cheating and their actor/targets.) Roy et al.
-(2021).
+fairness/cheating and their actor/targets.) roy-etal-2021-identifying.
 Participants.
@@
 In the relational system experiments, participants were tasked with selecting example documents for each identified theme, as well as determining concept relations for them.
-Following Pacheco et al.
-(2023), the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame).
+Following pacheco-2023-interactive, the supporting concepts considered were vaccination stances and morality frames (e.g., the identified theme “natural immunity” has an “anti-vax” stance, and is tied to the “purity” frame).
 Once participants were satisfied with their themes and selections, the system automatically coded the rest of the dataset.
@@
 We use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings.
-Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al.
-(2021); Pacheco et al.
-(2023).
+Our evaluation framework is comprised of three dimensions; consistency, cohesiveness & distinctiveness, and correctness, each of which uses metrics that are well-established in the literature NIPS2008_beed1360; 10.5555/3540261.3540416; pacheco-2023-interactive.
 Figure 2: Two sets of annotators use a particular HiTL system to find themes.
@@
 Consistency.
-Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (Cornish et al., 2014).
+Coders always run the risk of overgeneralizing or missing key themes, resulting in unclear and unsystematic results (cornish-2014-sage-cqa).
 We propose consistency as one dimension for measuring systematicity.
@@
 We want to account for the possibility of themes containing a different assortment of documents that otherwise remain relatively semantically similar.
-To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them Reimers and Gurevych (2019).
+To do this, we also measure the semantic similarity between themes using the S-BERT embeddings of the documents assigned to them reimers-gurevych-2019-sentence.
 We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme.
@@
 Annotators noted a lack of control and agency in the topic modeling system.
-Annotators expressed loss of control and agency when using the topic modeling system Fang et al.
-(2023) over other systems.
+Annotators expressed loss of control and agency when using the topic modeling system fang-2023-user over other systems.
 One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups".
@@
 In contrast to the clustering approach used by the other two systems, which appeared to yield cleaner partitions.
-One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al.
-(2023) model groups together despite the stark difference in message between the two.".
+One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the fang-2023-user model groups together despite the stark difference in message between the two.".
 Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended.
@@
 Despite the recent advances in the reasoning capabilities of LLMs, we observe that their performance at scale is not effective enough to be used for the purpose of propagating annotations.
-Additionally, running LLMs for the $5$ annotation configurations came with significant time and resource costs.
+Additionally, running LLMs for the 55 annotation configurations came with significant time and resource costs.
 This opens an opportunity for NLP researchers to make LLMs more reliable inductive reasoners, and to come up with prompting strategies that can allow LLMs to reliably classify documents in bulk, especially when working at scale.
@@
 We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output.
-This is particularly true for systems like Pacheco et al.
-(2023), where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding.
+This is particularly true for systems like pacheco-2023-interactive, where users have a wider bandwidth of interaction that can benefit from group consensus, and provide strong inductive biases for deductive coding.
 We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and accurate code assignments.
@@
 A larger study incorporating more tools could yield additional insights.
-(2) As prior work has shown Hoyle et al.
-(2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling.
+(2) As prior work has shown 10.5555/3540261.3540416, automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling.
 Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope.
@@
 Appendix A Topic Model Experimental Settings
-To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al.
-(2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$.
+To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as fang-etal-2021-query, which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$.
 The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability.
@@
 Finding Partitions Experts can find partitions in the space of unassigned instances.
-We currently support the K-means Jin and Han (2010) and Hierarchical Density-Based Clustering Mendelsohn et al.
-(2021) algorithms.
+We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mendelsohn-etal-2021-modeling algorithms.
 Text-based Queries Experts can type any query in natural language and find instances that are close to the query in the embedding space.
@@
 To do this, we aggregate all instances in the database.
-Explanations include theme distribution, coverage statistics, and t-sne plots van der Maaten and Hinton (2008).
+Explanations include theme distribution, coverage statistics, and t-sne plots JMLR:v9:vandermaaten08a.
 (a) Exploratory Operations
@@
 (b) Intervention Operations
-Table 3: Interactive Operations for the Pacheco et al. (2023) System
+Table 3: Interactive Operations for the pacheco-2023-interactive System
 Appendix C Semi-Structured Interview
@@
 The original work uses existing codes from a theoretical framework whereas we use codes defined by the user during this step, but this does not affect the overall process.
-code-book Gwet’s $AC_{1}$ Unlabeled Docs
+code-book Gwet’s $AC_{1}$ # Unlabeled Docs
 Sync 1 $0.42$ $5,548(6.5\%)$
@@
 Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work.
-We used the Llama 3.2 3B-Instruct (Grattafiori et al., 2024) model for all generation tasks:
+We used the Llama 3.2 3B-Instruct (grattafiori2024llama3herdmodels) model for all generation tasks:
 D.1 LLM Hyperparameters